# Photometric Inverse Rendering: Shading Cues Modeling and Surface Reflectance Regularization

### v1

@@ 1.1 Network Architectures
$f_{\Theta_{d}}(\boldsymbol{x},\boldsymbol{n},\boldsymbol{n},\boldsymbol{f})=\rho_{d}$. We use an 8-layer MLP featuring a hidden dimension of 256 and a skip connection at the fourth layer. The network inputs include the 3D coordinate $\boldsymbol{x}$ encoded with 10 frequencies, surface normal, and geometric feature. It outputs the diffuse albedo for point $\boldsymbol{x}$.
@@
$\gamma\left(\left\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\|,\boldsymbol{w}_{i}\cdot\boldsymbol{n},\rho_{r}\right)=\gamma$. We use a 4-layer MLP with width 128. The the dot product of normal and view direction uses 6 frequencies.
@@ 1.3 Inter-reflection Computation
To model the indirect illumination in scenes dynamically captured with directional lighting, we introduce an online computation approach that combines a differentiable layer and an importance sampling strategy. For a point $\boldsymbol{x}$ and view direction $\boldsymbol{w}_{i}$, we consider a single light bounce and employ ray marching towards the reflective direction:
@@
Given that the blending coefficient is conditioned on the roughness property of the 3D point, there exists a correlation between the roughness property and the blending coefficient, introducing additional ambiguity in material estimation. In our experiments, we found that detaching the roughness of the secondary intersection point $\boldsymbol{x}^{\prime}$ prior to its input into the blending coefficient network leads to a more precise material decomposition. Moreover, the process of gradient backpropagation starts from the image loss, through the residual component and into the material networks of the secondary intersection point $\boldsymbol{x}^{\prime}$, fosters the alignment of secondary point radiance with inter-reflection cues. In our experiments, we discovered that disabling the optimization of local geometry at the secondary point reduces the complexity of the optimization process, leading to improved geometric reconstruction, especially in concave areas.
@@ 1.5 Training Details
The training process requires approximately 9 hours on a single RTX3090 GPU with 24GB of memory. We start by training NeuS over 100,000 iterations to initialize the geometry and diffuse albedo networks. For each training iteration, we utilize 512 randomly sampled pixels, employing an $\ell_{1}$ loss along with an eikonal regularization loss. Prior to the rendering phase, we derive the feature maps of images by the pre-trained ViT-S/8 model [9] and executed 10,000 iterations with $\lambda_{4}$ set to $1.0$ to extract the DINO feature from 2D feature maps to 3D surfaces. During the physics-based surface rendering stage with total 50,000 iterations, we fixed the geometry and lighting to warmup the BRDFs network for 2,000 iterations to stablize the process, and subsequently, we carried out a joint optimization of the lighting, geometry and BRDFs. The training of blending coefficient network started at the 10,000th iteration. We set the size of rendered image patch as $128\times 128$ and loss weights to $\lambda_{1}=10^{-4}$, $\lambda_{2}=0.1$, $\lambda_{3}=10^{-5}$ and $\lambda_{4}=10^{-5}$. All networks are optimized by corresponding Adam optimizers with learning rate $10^{-4}$.
@@ 2 More Details for the Dataset
For capturing real-world images, we employed an iPhone 15 to shoot in RAW format, ensuring a linear camera response. Across all photos, we maintained consistent settings for the camera’s exposure time, focus, and white balance. Specifically, the ISO value and shutter speed (exposure time) were fixed at 100 and 1/250s, respectively, with the white balance adjusted to 3,800 Kelvin degrees. Our collection encompasses 5 scenes: Toy, fruit, Panda, Assassin and Bear, with the number of images per scene varying from 120 to 400. Camera poses were derived using COLMAP [50], and objects were scaled to fit within a unit sphere based on the reconstructed point cloud. The photography sessions took place in a darkroom, positioning the camera 0.15 to 0.3 meters away from each object. To achieve comprehensive coverage, we systematically moved the camera in a spiral pattern around the subjects. The separation between the camera lens and the flashlight on the iPhone, roughly 0.015m, results in an approximate 3-degree variation between viewing and lighting angles at a standard distance of 0.25m from the object.
@@ 3.1 Results on Synthetic Data
In Table 5, we offer comprehensive results for each synthetic scene captured under casual conditions. Additionally, we provide a qualitative comparison of novel view rendering and material decomposition between our method and earlier methods, as illustrated in Fig. 11. For the dragon scene, our method produces a diffuse albedo with less indirect illumination incorporated into the materials. In the horse scene, our material decomposition results demonstrate a reduced influence of self-shadows, showing a closer alignment with the ground truth than those obtained with IRON. Even in extreme concave regions, our method is more robust than previous method as shown in Fig. 12.
@@ 3.2 Results on Real Data
In Fig. 14, we present our dataset’s novel view rendering and material decomposition outcomes. The IRON method often incorporates indirect illumination into the diffuse albedo, particularly in concave regions, as observed in the Fruit scene. Additionally, specular albedoes produced by IRON method are adversely affected by self-shadows and inter-reflections, as highlighted in specific boxes.
In Table 5, we provide a quantitative comparison that underscores the enhanced performance of our method compared to IRON in terms of novel view rendering within the DRV real dataset. Figure 15 and Fig. 16 complement this with side-by-side qualitative comparisons of our method against IRON regarding to material decomposition. Leveraging DINO regularization for surface decomposition, which effectively clusters similar materials, our approach produces more accurate results for material decomposition, especially in scenarios with a skewed view distribution. We observe that IRON’s evaluation metrics for the Dragon scene slightly exceed those of our method, this disparity is primarily due to its collocated camera-lighting setup, which inherently minimizes the occurrence of self-shadows within the scene.

### v2

@@ 1.1 Network Architectures
$f_{\Theta_{d}}(\boldsymbol{x},\boldsymbol{n},\boldsymbol{n},\boldsymbol{f})=\rho_{d}$. We use an 8-layer MLP featuring a hidden dimension of 256 and a skip connection at the fourth layer. The network inputs include the 3D coordinate $\boldsymbol{x}$ encoded with 10 frequencies, surface normal, and geometric features. It outputs the diffuse albedo for point $\boldsymbol{x}$.
@@
$\gamma\left(\left\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\|,\boldsymbol{w}_{i}\cdot\boldsymbol{n},\rho_{r}\right)=\gamma$. We use a 4-layer MLP with a width of 128. The dot product of normal and view direction uses 6 frequencies.
@@ 1.3 Inter-reflection Computation
To model the indirect illumination in scenes dynamically captured with directional lighting, we introduce an online computation approach that combines a differentiable layer and an importance sampling strategy. For a point $\boldsymbol{x}$ and view direction $\boldsymbol{w}_{i}$, we consider a single light bounce and employ a ray marching towards the reflective direction:
@@
Given that the blending coefficient is conditioned on the roughness property of the 3D point, there exists a correlation between the roughness property and the blending coefficient, introducing additional ambiguity in material estimation. In our experiments, we found that detaching the roughness of the secondary intersection point $\boldsymbol{x}^{\prime}$ prior to its input into the blending coefficient network leads to a more precise material decomposition. Moreover, the process of gradient backpropagation starts from the image loss, through the residual component, and into the material networks of the secondary intersection point $\boldsymbol{x}^{\prime}$, fostering the alignment of secondary point radiance with inter-reflection cues. In our experiments, we discovered that disabling the optimization of local geometry at the secondary point reduces the complexity of the optimization process, leading to improved geometric reconstruction, especially in concave areas.
@@ 1.5 Training Details
The training process requires approximately 9 hours on a single RTX3090 GPU with 24GB of memory. We start by training NeuS over 100,000 iterations to initialize the geometry and diffuse albedo networks. For each training iteration, we utilize 512 randomly sampled pixels, employing an $\ell_{1}$ loss along with an eikonal regularization loss. Prior to the rendering phase, we derive the feature maps of images by the pre-trained ViT-S/8 model [9] and executed 10,000 iterations with $\lambda_{4}$ set to $1.0$ to extract the DINO feature from 2D feature maps to 3D surfaces. During the physics-based surface rendering stage with a total of 50,000 iterations, we fixed the geometry and lighting to warm up the BRDFs network for 2,000 iterations to stabilize the process, and subsequently, we carried out a joint optimization of the lighting, geometry, and BRDFs. The training of the blending coefficient network started at the 10,000th iteration. We set the size of rendered image patch as $128\times 128$ and loss weights to $\lambda_{1}=10^{-4}$, $\lambda_{2}=0.1$, $\lambda_{3}=10^{-5}$ and $\lambda_{4}=10^{-5}$. All networks are optimized by corresponding Adam optimizers with learning rate $10^{-4}$.
@@ 2 More Details for the Dataset
For capturing real-world images, we employed an iPhone 15 to shoot in RAW format, ensuring a linear camera response. Across all photos, we maintained consistent settings for the camera’s exposure time, focus, and white balance. Specifically, the ISO value and shutter speed (exposure time) were fixed at 100 and 1/250s, respectively, with the white balance adjusted to 3,800 Kelvin degrees. Our collection encompasses 5 scenes: Toy, fruit, Panda, Assassin, and Bear, with the number of images per scene varying from 120 to 400. Camera poses were derived using COLMAP [50], and objects were scaled to fit within a unit sphere based on the reconstructed point cloud. The photography sessions took place in a darkroom, positioning the camera 0.15 to 0.3 meters away from each object. To achieve comprehensive coverage, we systematically moved the camera in a spiral pattern around the subjects. The separation between the camera lens and the flashlight on the iPhone, roughly 0.015m, results in an approximate 3-degree variation between viewing and lighting angles at a standard distance of 0.25m from the object.
@@ 3.1 Results on Synthetic Data
In Table 5, we offer comprehensive results for each synthetic scene captured under casual conditions. Additionally, we provide a qualitative comparison of novel view rendering and material decomposition between our method and earlier methods, as illustrated in Fig. 11. For the dragon scene, our method produces a diffuse albedo with less indirect illumination incorporated into the materials. In the horse scene, our material decomposition results demonstrate a reduced influence of self-shadows, showing a closer alignment with the ground truth than those obtained with IRON. Even in extremely concave regions, our method is more robust than the previous method, as shown in Fig. 12.
@@ 3.2 Results on Real Data
In Fig. 14, we present our dataset’s novel view rendering and material decomposition outcomes. The IRON method often incorporates indirect illumination into the diffuse albedo, particularly in concave regions, as observed in the Fruit scene. Additionally, specular albedoes produced by the IRON method are adversely affected by self-shadows and inter-reflections, as highlighted in specific boxes.
In Table 5, we provide a quantitative comparison that underscores the enhanced performance of our method compared to IRON in terms of novel view rendering within the DRV real dataset. Figure 15 and Fig. 16 complement this with side-by-side qualitative comparisons of our method against IRON regarding material decomposition. Leveraging DINO regularization for surface decomposition, which effectively clusters similar materials, our approach produces more accurate results for material decomposition, especially in scenarios with a skewed view distribution. We observe that IRON’s evaluation metrics for the Dragon scene slightly exceed those of our method, this disparity is primarily due to its collocated camera-lighting setup, which inherently minimizes the occurrence of self-shadows within the scene.